# Redundancy of the cosmological evolution equations and its relationship with the initial conditions

### v2

@@ Abstract
It is known that in Friedmann-Lemaitre-Robertson-Walker cosmology one has more number of dynamical equations, compared to the number of unknown variables. This fact makes some equations redundant. The situation becomes complicated because all the relevant differential equations in cosmology are not of the same order. In this article we study the fate of the redundant equations. We show that this redundancy is inevitable in general relativity and a modern modification of gravity, called $f(R)$ gravity. It is shown that this redundancy predicts a special role to one of the Friedmann equations, which constrains the initial values of the problem. It is shown that the initial constraint condition itself evolves with time. Our method of analyzing the dynamical structure of the theories relies on an operational approach and can be generalized further.
@@ 1 Introduction
The basics of modern cosmology are taught at the masters level in various institutions. The standard paradigm of introductory cosmology is well addressed in Refs. [1, 2, 3] whereas Refs. [4, 5, 6] focusses more on general relativity (GR). These books give a general understanding based on which someone can build up his/her conceptions in cosmology. Readers who will like to know more about the various aspects of modern cosmology based on the Friedmann-Lemaitre-Robertson-Walker (FLRW) spacetime can consult the excellent pedagogical articles in Refs. [7, 8, 9, 10, 11]. While going through the basics of cosmology one first learns to tackle the dynamical equations in cosmology, commonly called the Friedmann equations. These Friedmann equations are nonlinear ordinary differential equations which originate from the Einstein equations in GR which are second order nonlinear partial differential equations in three spatial and one time dimensions. While dealing with these equations one first observes that even in the simplest situation, where the universe only contains one barotropic fluid, there is a problem of redundancy. There are simply more number of cosmological evolution equations than required. The situation is more challenging as all the available differential equations are not of the same order. Naturally the question arises why such a problem appears at all and how are we going to obtain proper cosmological development from these equations in a consistent manner. In this article we will primarily be addressing this redundancy problem in cosmology. We will first show that this redundancy problem is intimately related to the initial value problem in cosmology. It turns out that any arbitrary initial condition will not produce proper cosmological dynamics, only initial conditions satisfying a particular constraint will produce proper dynamics. Then we show that the redundancy problem in cosmology has its own logic in GR. We will also briefly state how our analysis in GR can be generalized in a more modern variation of a theory of gravity, called $f(R)$ gravity.
The article is intended for readers who have taken a basic course of introductory cosmology and knows how to solve the Friedmann equations in standard scenarios: As in matter domination or in radiation domination [^1] We assume that the intended reader to be aware of at least some of the references cited in the last paragraph.. What has been presented in this article has deeper connections to the theory of second order nonlinear partial differential equations, involving time development, in three dimensional space.
The interesting point which becomes apparent from the present article is related to the fact that we can learn many interesting properties related to cosmological development in GR, as well as in $f(R)$ gravity, from an operational approach where our aim is solely to solve the dynamical equations unambiguously. Such an approach ultimately leads to a thorough understanding of the logical structure of the theories.
@@ 2 A single barotropic fluid in FLRW spacetime and the Einstein equation
The spacetime metric, in cosmology, is the isotropic and homogeneous FLRW spacetime:
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right]\,$, (1)
where $a(t)$ is the scale-factor for the expanding or contracting spacetime and $k$ is a constant which can take values $\pm 1$ or 0. If $k=0$, the curvature of the spatial hypersurfaces in FLRW spacetime is zero, indicating flat spatial hypersurfaces. On the other hand k=+1,1k=+1,\,-1, specifies that the spatial hypersurfaces are positively curved (as on the surface of a sphere) or negatively curved (as in a saddle), respectively. The spatial coordinate $r$, used in the above line element is called a comoving coordinate. One must be careful in interpreting the above line element. We are using units where speed of light is one. As a consequence length and time have the same dimension. We can interpret $ds^{2}$ having a dimension of squared length. In this unit what will be the dimensions of $r$, $a(t)$ and $k$? When $k=0$, $r$ does have the dimension of length and $a(t)$ can be dimensionless. When $k=\pm 1$ and dimensionless, $r$ cannot carry length dimension. In such cases $a(t)$ carries the dimension of length. As $k$ comes out from the geometry of the FLRW spacetime, we will like to keep it dimensionless.
In this work we will assume that the FLRW spacetime is seeded by a single component barotropic fluid whose equation of state (EoS):
$P=\omega\rho\,$, (2)
where $\omega$, which is a dimensionless quantity in our system of units, is often called the EoS parameter. Here $\rho$ is the energy-density of the fluid and $P$ specifies its isotropic pressure. The EoS parameter is a constant. In presence of a barotropic fluid, the Einstein equations of FLRW cosmology are:
@@
where the dot specifies a derivative with respect to time, $t$. The first of the above two equations is called the second Friedmann equation by many authors. Subtracting Eq. (3) from Eq. (4) one gets
@@
Many authors call the above equation to be the first Friedmann equation. The last equation is also called the acceleration equation. In this work we will call Eq. (3) as the Friedmann equation, the second equation, Eq. (4), will be called as the pressure equation and the last equation will be named as the acceleration equation.
Except the above three equations, which govern the dynamics of the system, there is another one. This equation comes from the energy-momentum conservation condition. The energy-momentum conservation condition in GR is an outcome of the Bianchi identity. In FLRW spacetime this condition becomes
@@
Now we have four dynamical equations for the development of the FLRW spacetime. We emphasize here that, although we know that we have two independent differential equations coming out from the Einstein equation (the Friedmann equation and the the pressure equation), they cannot be simultaneously solved. The reason being, none of them contain derivatives of $\rho$ or $P$. In such a case one is bound to work with the continuity equation as it is the only one which has a time derivative of $\rho$. We have three functions to determine: $a(t)$, $\rho(t)$ and $P(t)$. Using the EoS we know $P(t)$ can be expressed in terms of the energy density and hence we have only two unknown functions. How do we figure out the development of this dynamical system?
We can now be more specific about the problem of time evolution in the FLRW system. Is it always possible that the four differential equations agree during $t_{i}\leq t\leq t_{f}$ while the actual dynamics is produced by only two differential equations satisfying some specific initial conditions, keeping $a(t)>0$ and $\rho(t)>0$ throughout the time interval $t_{i}\,,\,t_{f}]$? The answer turns out to be positive for a certain set of initial conditions. For initial conditions which are outside the preferred set, all the four equations will never match and the system will behave as an overdetermined system. In the last case we have to conclude that the Einstein equation does not yield any specific solution. We present various cases to illustrate the various possibilities in the present problem.
To present our arguments we will first work with cosmologies involving purely expanding or contracting phases. Later we will discuss the cases where an expanding (contracting) phase can convert to a contracting (expanding) phase.
@@ 3 The various cases in an ever expanding or ever contracting universe
In this section we will discuss about purely expanding or contracting phases in cosmology. Throughout the relevant interval, $t_{i}\leq t\leq t_{f}$, we assume that $\dot{a}\neq 0$. We will relax the last assumption in the next section and show that for intervals where we have $\dot{a}(t_{0})=0$, at some point $t_{0}$ in the interval, we can still obtain specific solutions of the Einstein equation. We will first try to see whether we can reproduce the differential equations left out from our chosen pair of equations. Although the four differential equations are produced essentially from two independent dynamical equations, it will be seen that one may not get back all the equations from any pair of relevant equations. We will then try to see whether the solution produced from the chosen pair of equations simultaneously satisfy the other equations.
Before we start we must point out that the time interval $t_{i}\,,\,t_{f}]$ has to be chosen carefully. We have demanded the interval to be such that the scale-factor always increases or decreases with time inside the interval. We have tacitly assumed that $a(t)>0$ inside the interval. The last condition may not be true for any time interval. For a contracting spacetime, the dynamics can produce $a(t_{s})=0$ for some $t_{i}\leq t_{s}\leq t_{f}$. In such cases the system has reached a singularity. In this case one has to redefine the interval in such a way that the upper limit of the relevant interval, $t_{f}$, always remains smaller than $t_{s}$. In other words, our interval must be chosen in such a way that there are no singularities inside this interval.
@@ 3.1 Choosing the Friedmann equation and the continuity equation
When we choose the Friedmann equation and the continuity equation, we see that both of them are first order differential equations. They require two independent initial conditions: $a(t_{i})$ and $\rho(t_{i})$. We have not chosen the pressure equation and the acceleration equation in our scheme. To obtain the value of $\ddot{a}$ we first take a time derivative of both the sides of the Friedmann equation and get
@@
and then utilizing the Friedmann equation, to replace $8\pi G\rho$, we get the pressure equation. This shows that the values of $a(t)$ and $\rho(t)$ generated from the Friedmann equation also satisfies the pressure equation for any $t$ inside the interval $t_{i}\,,\,t_{f}]$ for cases where $\dot{a}\neq 0$ inside the time interval. Now, one can again combine the pressure equation with the Friedmann equation to obtain the acceleration equation. As a consequence of this the generated values of $a(t)$ and $\rho(t)$ will also satisfy the acceleration equation. Hence all the relevant dynamical equations are satisfied simultaneously in this case.
@@ 3.2 Choosing the pressure equation and the continuity equation
This equation shows that the Friedmann equation may not hold when we obtain the dynamics of the system from the pressure equation and the continuity equation. If and only if $F(t)=0$ for all instants in the interval $t_{i}\,,\,t_{f}]$, the Friedmann equation will be valid throughout the interval. In general for arbitrary initial conditions, $F(t)$ is never zero at any point inside the interval. Consequently for arbitrary initial conditions this dynamical evolution will not satisfy the Friedmann equation. The resultant dynamical evolution cannot be a valid solution of the cosmological evolution equations.
@@ 3.4 The link between all the choices is in the initial condition
From our previous discussion we have seen that all the pairs of chosen differential equations may not reproduce the other differential equations. All the previous discussion was based on arbitrary choice of initial conditions. We did not mention any scheme for comparing initial conditions. The pressure equation and continuity equation as well as the acceleration equation and the continuity equation requires three numbers in the initial condition: $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$. There is no unique way to choose this triplet of numbers. On the other hand the Friedmann equation and the continuity equation only requires two initial conditions: $a(t_{i})$ and $\rho(t_{i})$. Apparently it seems it is impossible to specify the initial condition which can be applied to all the three pairs of equations. If we treat all the differential equations as purely dynamical equations then it is really impossible to find out an equivalent initial condition.
The above problem can be tackled if we reinterpret the Friedmann equation as a constraint equation at the initial point $t_{i}$. The Friedmann equation is the only equation which can be interpreted as an algebraic equation connecting the values of $\dot{a}(t_{i})$, $a(t_{i})$ and $\rho(t_{i})$ for some values of $k$ and $\Lambda$. Once we choose the initial conditions from the Friedmann equation itself, using it as a constraint equation, we see that the Friedmann equation remains valid throughout the time interval $t_{i}\,,\,t_{f}]$. This happens because of the nature of Eq. (14). If the initial condition satisfies the Friedmann constraint then $F(t_{i})=0$ and consequently $\dot{F}(t_{i})=0$. As we chose the pressure equation and the continuity equation to generate the dynamics, we see that if the initial condition is chosen from the Friedmann constraint then the Friedmann equation remains valid throughout the time interval. The generated dynamics is naturally a solution of the Friedmann equation.
Next we focus our attention to the dynamics generated by the acceleration equation and the continuity equation. In this case we do not know the value of $\tilde{k}$. If the initial condition is chosen from the Friedmann equation, with a specific value of $k$, then what happens? We see from Eq. (20), this initial condition can only be chosen when $\tilde{k}=k$. In this case, the Friedmann constraint gives the information about the spatial curvature of the initial 3-dimensional hypersurface. Not only the initial condition fixes the value of $k$, now $\tilde{F}(t)=F(t)$, where $F(t)$ is defined as in Eq. (13). In this case also we see, from Eq. (20), that if $\tilde{F}(t_{i})=0$ then $\tilde{F}(t)=0$, at all instants in the interval $t_{i}\,,\,t_{f}]$. All the four relevant dynamical differential equations will have the same solution if we choose the initial condition via the Friedmann constraint.
@@ 4.2 When $t_{i}<t_{0}<t_{f}$
when $\dot{a}(t=0)=0$. We will use these properties and shift the origin of time to $t=t_{0}$ so that what used to be called $t_{0}$ becomes the origin of time $t=0$. The initial time interval now becomes $t_{i}-t_{0}\,,\,t_{f}-t_{0}]$. In shifted time we have $\dot{a}(0)=0$ which implies a symmetric gravitational collapse. The initial conditions can be well specified in the shifted time as $\rho(0)$ and $a(0)$. Now using any pair of differential equations we can generate the solution of the Einstein equation in the expanding phase, in the interval $t_{i}-t_{0},0]$. In this interval all the solutions will agree if all of them satisfy the Friedmann constraint initially. The symmetry of the Einstein equations now guarantee that these solutions must also agree on the contracting phase. If one now wants to go back to the old time interval, one has to just shift the time origin back to $t_{i}$.
@@ 5 On the logical consistency of the theory
From our previous discussion it is seen that one can model cosmological dynamics from any relevant pair of equations (presented in this article) and even if those equations do not involve the Friedmann equation specifically, one cannot forget about the Friedmann equation. All cosmological initial conditions must satisfy the Friedmann equation. Although we showed the relevance of this statement in a simple case, the last statement is valid in more general cases as will be shown later in this article.
@@
The results of this article are related to the initial value problem and its solution in cosmology. Can one set up proper initial conditions in GR and solve the Einstein equations consistently in the future, where one is using the FLRW spacetime? This is the initial value problem in cosmology and it is known that the answer is in the affirmative. A formal discussion on the initial value problem in cosmology and its solution can be found in Refs. [12, 13]. In the rest of this section we discuss about the logical consistency of cosmological dynamics in GR.
Till now what has been discussed indicates that the complexity of the situation is primarily related to the availability of both first order and second order differential equations. Naturally the question arises: Is this a peculiarity of the cosmological equations in GR? In this section we will show that the dynamical structure of the cosmological theory, as predicted by GR, is logically consistent and makes perfect sense.
@@
To understand the logical consistency in the present case, we will like to specify here that some form of a first order differential equation is inevitable in the present case if we have to make the theory meaningful. If we leave out quantum effects, we may safely assume that there cannot be matter creation from the gravitational field. In such a case we must have some condition which predicts that the evolution equation of energy density is a first order differential equation and its form must be $\dot{\rho}\propto\rho$. This is a very general requirement, as for such a relation if $\rho$ is zero initially then $\dot{\rho}$ is also zero initially and hence the matter energy density must remain zero for all times. A second order differential equation in $\rho(t)$ will never satisfy the general requirement, as in this case even if initially $\rho$ and $\dot{\rho}$ is zero, still $\rho(t)$ can attain finite non-zero values in the future.

### v3

@@ Abstract
It is known that in Friedmann-Lemaitre-Robertson-Walker cosmology one has more number of dynamical equations, compared to the number of unknown variables. This fact makes some equations redundant. The situation becomes complicated because all the relevant differential equations in cosmology are not of the same order. In this article we study the fate of the redundant equations. We show that this redundancy is inevitable in general relativity. It is shown that this redundancy is primarily responsible for a special role of one of the Friedmann equations, which constrains the initial values of the problem. Our method of analyzing the dynamical structure of the theories relies on an operational approach and can be generalized further.
@@ 1 Introduction
While going through the basics of cosmology one first learns to tackle the dynamical equations, commonly called the Friedmann equations in Friedmann-Lemaitre-Robertson-Walker (FLRW) spacetime  [^1] The standard paradigm of introductory cosmology is well addressed in Refs. [1, 2, 3] whereas Refs. [4, 5, 6] focusses more on general relativity. Some excellent pedagogical articles on introductory cosmology can also be found in Refs. [7, 8, 9, 10, 11].. These Friedmann equations are nonlinear ordinary differential equations which originate from the Einstein equations in general relativity (GR). While dealing with these equations one first observes that even in the simplest situation, where the universe only contains one barotropic fluid, there is a problem of redundancy. There are simply more number of cosmological evolution equations than required. The situation is more challenging as all the available differential equations are not of the same order. Naturally the question arises why such a problem appears at all and how are we going to obtain proper cosmological development from these equations in a consistent manner. In this article we will primarily be addressing this redundancy problem in cosmology. We will first show that this redundancy problem is intimately related to the initial value problem in cosmology. It turns out that any arbitrary initial condition will not produce proper cosmological dynamics, only initial conditions satisfying a particular constraint will produce proper dynamics. Then we show that the redundancy problem in cosmology has its own logic in GR.
The interesting point which becomes apparent from the present article is related to the fact that we can learn many important properties related to cosmological development from an operational approach, where our aim is solely to solve the dynamical equations unambiguously. Such an approach ultimately leads to a thorough understanding of the logical structure of the theory.
@@ 2 A single barotropic fluid in FLRW spacetime and the Einstein equation
In this work we will assume that the FLRW spacetime is seeded by a single component barotropic fluid whose equation of state (EoS): $P=\omega\rho\,$, where $\omega$, is a dimensionless quantity in our system of units (where $c=1$). Here $\rho$ is the energy-density of the fluid and $P$ specifies its isotropic pressure. The EoS parameter is a constant. In presence of a barotropic fluid and cosmological constant $\Lambda$, the Einstein equations of FLRW spacetime, with a scale-factor $a(t)$, are given as:
@@
where the dot specifies a derivative with respect to time, $t$, and $k$ is a constant which can take values $\pm 1$ or 0. If $k=0$, the curvature of the spatial hypersurfaces in FLRW spacetime is zero, on the other hand k=+1,1k=+1,\,-1, specifies that the spatial hypersurfaces are positively curved (as on the surface of a sphere) or negatively curved (as in a saddle), respectively.
Subtracting Eq. (1) from Eq. (2) one gets another dynamical equation in FLRW cosmology:
@@
In this work we will call Eq. (1) as the Friedmann equation, the second equation, Eq. (2), will be called as the pressure equation and the last equation will be named as the acceleration equation.
Except the above three equations we have the energy-momentum conservation condition in GR, which is an outcome of the Bianchi identity. In FLRW spacetime this condition becomes
@@
There are now four dynamical equations for the development of the FLRW spacetime. It must be noted that, although we have two independent differential equations coming out from the Einstein field equation (the Friedmann equation and the the pressure equation), they cannot be simultaneously solved as none of them contain derivatives of $\rho$ or $P$. In such a case one is bound to work with the continuity equation as it is the only one which has a time derivative of $\rho$. There are three functions to be determined: $a(t)$, $\rho(t)$ and $P(t)$. Using the EoS, $P(t)$ can be expressed in terms of the energy density and hence, there are only two unknown functions.
Is it always possible that the four differential equations agree during $t_{i}\leq t\leq t_{f}$ while the actual dynamics is produced by only two differential equations satisfying some specific initial conditions, keeping $a(t)>0$ and $\rho(t)>0$ throughout the time interval $t_{i}\,,\,t_{f}]$? The answer turns out to be positive for a certain set of initial conditions. For initial conditions which are outside the preferred set, all the four equations will never match and the system will behave as an overdetermined system. To present our arguments we will first work with cosmologies involving purely expanding or contracting phases. Later we will discuss the cases where an expanding (contracting) phase can convert to a contracting (expanding) phase.
@@ 3 The various cases in an ever expanding or ever contracting universe
In this section we will discuss about purely expanding or contracting phases in cosmology. Throughout the relevant interval, $t_{i}\leq t\leq t_{f}$, we assume that $\dot{a}\neq 0$. It is natural to assume that $a(t)>0$ inside the interval. The last condition may not be true for any time interval. For a contracting spacetime, the dynamics can produce $a(t_{s})=0$ for some $t_{i}\leq t_{s}\leq t_{f}$. In such cases the system has reached a singularity. In this case one has to redefine the interval in such a way that the upper limit of the relevant interval, $t_{f}$, always remains smaller than $t_{s}$. In other words, our interval must be chosen in such a way that there are no singularities inside this interval.
@@ 3.1 Choosing the Friedmann equation and the continuity equation
When we choose the Friedmann equation and the continuity equation, we see that both of them are first order differential equations. They require two independent initial conditions: $a(t_{i})$ and $\rho(t_{i})$. We have not chosen the pressure equation and the acceleration equation in our scheme. Are those equations consistent with the dynamics produced from the two chosen equations? To answer that we proceed as follows.
To obtain the value of $\ddot{a}$ we first take a time derivative of both the sides of the Friedmann equation and get
@@
and then utilizing the Friedmann equation, to replace $8\pi G\rho$, we get the pressure equation. This shows that the values of $a(t)$ and $\rho(t)$ generated from the Friedmann equation also satisfies the pressure equation for any $t$ inside the interval $t_{i}\,,\,t_{f}]$ for cases where $\dot{a}\neq 0$ inside the time interval. One can again combine the pressure equation with the Friedmann equation to obtain the acceleration equation. As a consequence of this the generated values of $a(t)$ and $\rho(t)$ will also satisfy the acceleration equation. Hence all the relevant dynamical equations are satisfied simultaneously in this case.
@@ 3.2 Choosing the pressure equation and the continuity equation
This equation shows that the Friedmann equation may not hold when we obtain the dynamics of the system from the pressure equation and the continuity equation. If and only if $F(t)=0$ for all instants in the interval $t_{i}\,,\,t_{f}]$, the Friedmann equation will be valid throughout the interval. In general for arbitrary initial conditions, $F(t)$ is never zero for all points inside the time interval. Consequently, for arbitrary initial conditions this dynamical evolution will not satisfy the Friedmann equation and the resultant dynamical evolution may not be a valid solution of the cosmological evolution equations.
@@ 3.4 The link between all the choices is in the initial condition
The pressure equation and continuity equation, as well as, the acceleration equation and the continuity equation, requires three numbers in the initial condition: $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$. On the other hand the Friedmann equation and the continuity equation only requires two initial conditions: $a(t_{i})$ and $\rho(t_{i})$. If we treat all the differential equations as purely dynamical equations then it is really impossible to find out an equivalent initial condition.
The above problem can be tackled if we reinterpret the Friedmann equation as a constraint equation at the initial point $t_{i}$. The Friedmann equation is the only equation which can be interpreted as an algebraic equation connecting the values of $\dot{a}(t_{i})$, $a(t_{i})$ and $\rho(t_{i})$ for some values of $k$ and $\Lambda$. Once we choose the initial conditions from the Friedmann equation itself, using it as a constraint equation, we see that the Friedmann equation remains valid throughout the time interval $t_{i}\,,\,t_{f}]$. This happens because of the nature of Eq. (12). If the initial condition satisfies the Friedmann constraint then $F(t_{i})=0$ and consequently $\dot{F}(t_{i})=0$. As we chose the pressure equation and the continuity equation to generate the dynamics, we see that if the initial condition is chosen from the Friedmann constraint then the Friedmann equation remains valid throughout the time interval.
Next we focus our attention to the dynamics generated by the acceleration equation and the continuity equation. If the initial condition is chosen from the Friedmann equation, with a specific value of $k$, then Eq. (18) predicts $\tilde{k}=k$. The Friedmann constraint gives the information about the spatial curvature of the initial 3-dimensional hypersurface. Not only the initial condition fixes the value of $k$, now $\tilde{F}(t)=F(t)$, where $F(t)$ is defined as in Eq. (11). In this case also we see, from Eq. (18), that if $\tilde{F}(t_{i})=0$ then $\tilde{F}(t)=0$, at all instants in the interval $t_{i}\,,\,t_{f}]$. All the four relevant dynamical differential equations will have the same solution if we choose the initial condition via the Friedmann constraint.
@@ 4.2 When $t_{i}<t_{0}<t_{f}$
when $\dot{a}(t=0)=0$. We will use these properties and shift the origin of time to $t_{0}$ so that what we previously called the time instant $t_{0}$ becomes the origin of time $t=0$. The initial time interval now becomes $t_{i}-t_{0}\,,\,t_{f}-t_{0}]$. In shifted time we have $\dot{a}(0)=0$ which implies a symmetric gravitational collapse. The initial conditions can be well specified in the shifted time as $\rho(0)$ and $a(0)$. Now using any pair of differential equations we can generate the solution of the Einstein equation in the expanding phase, in the interval $t_{i}-t_{0},0]$. In this interval all the solutions will agree if all of them satisfy the Friedmann constraint initially. The symmetry of the Einstein equations now guarantee that these solutions must also agree on the contracting phase. If one now wants to go back to the old time interval, one has to just shift the time origin back to $t_{i}$.
@@ 5 On the logical consistency of the theory
The results of this article are related to the initial value problem and its solution in cosmology. A formal discussion on the initial value problem in cosmology, using techniques of differential geometry, can be found in Refs. [12, 13]. In this article we focus on cosmological dynamics to infer upon the initial value problem in cosmology. In the rest of this section we discuss about the logical consistency of cosmological dynamics in GR.
@@
If we leave out quantum effects, we may safely assume that there cannot be matter creation from the gravitational field. In such a case we must have some condition which predicts $\rho(t)=0$ in the interval $t_{i}\,,\,t_{f}]$ if $\rho(t_{i})=0$. Such a relation is there in GR, as the theory is formulated in such a way that the desired condition is produced due to a tensor identity, called the (second) Bianchi identity. This identity predicts the conservation of matter energy-momentum tensor, which yields a first order differential equation: where $\dot{\rho}(t)\propto\rho(t)$. This is a very general requirement, as for such a relation if $\rho$ is zero initially then $\dot{\rho}$ is also zero initially and hence the matter energy density must remain zero for all times.